OpenStudy (anonymous):

What is the sum of the arithmetic sequence 8, 14, 20 …, if there are 24 terms? Help Please

OpenStudy (anonymous):

1,842 1,844 1,846 1,848

OpenStudy (anonymous):

^ possible answers

OpenStudy (anonymous):

a=8 d= 20-14 = 14-8 = 6 24th term=? a24(last term)=a+(n-1)d =8+(24-1)6 = 8+23*6 = 8+138 = 146 Sum of n terms = n/2 (a+l) l=last term =24/2 (8+146) = 12*154 =1848

OpenStudy (anonymous):

thanks